The sum of three numbers is 119. The first number is 5 more than the third. The second number is 4 times the third. What are the three numbers?
Answer by lwsshak3(11628)

The sum of three numbers is 119. The first number is 5 more than the third. The second number is 4 times the third. What are the three numbers?
***
let x=3rd number
x+5=1st number
4x=2nd number
x+x+5+4x=119
6x=114
x=19
4x=76
x+5=24